write an equation of the line that is perpendicular to 3x-4y=9 and contains (-4,-5)
Answer by Alan3354(69443)

write an equation of the line that is perpendicular to 3x-4y=9 and contains (-4,-5)
------------
Find the slope of the given line.
The slope of lines perpendicular is the negative inverse.
----
Use y-y1 = m*(x-x1) where m = slope and (x1,y1) is the point.